OpenStudy (gabylovesyou):

Select the equations that are parallel and perpendicular to y = 1/4x + 7 and that pass through the point (-8,9) parallel: y = 1/4x - 11 perpendicular: y = -4x + 23 Parallel: y = -4x + 11 Perpendicular: y = 1/4x - 4 Parallel: y = 4x - 7 perpendicular: y = -1/4x - 7 parallel: y = 1/4x + 11 perpendicular: y = -4x - 23

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What is the slope of the line y = 1/4x + 7

OpenStudy (gabylovesyou):

how do i find it?

OpenStudy (sirm3d):

the last one is the correct answer

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

the slope of y = mx+b is m

jimthompson5910 (jim_thompson5910):

in the case of y = 1/4x + 7, what is m

OpenStudy (gabylovesyou):

1/4

jimthompson5910 (jim_thompson5910):

so that's the slope of any line parallel to that one

jimthompson5910 (jim_thompson5910):

so that allows you to rule out choices B and C

OpenStudy (gabylovesyou):

is it D? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

it is, nice job